OpenStudy (anonymous):

What is the solution to the rational equation 4/3 + 2/(x-5) = -6/(3x - 15) ? a. x = –8 b. x = –2 c. x = 2 d. x = 8 I think its c, am i right?

OpenStudy (anonymous):

so what happens when you put your answer, x = 2, into the equation?

OpenStudy (anonymous):

LHS: 4/3 + 2/(2-5) = 4/3 + 2/-3 = 2/3 RHS: -6/(3(2) - 15) = -6/-9 = 2/3 so you are correct. do you see how you can check your answer by substituting back?

OpenStudy (anonymous):

so i was right..

OpenStudy (anonymous):

yes, and now you know how to check your answers yourself :)

OpenStudy (anonymous):

yep, thanks!